Circular No. 7130

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999au, 1999av, 1999aw, 1999ax, 1999ay Five new supernova discoveries have been reported — two by A. Gal-

Yam and D. Maoz (on behalf of the Wise Observatory Optical Transients Search, at the Wise 1-m reflector; cf. IAUC 7055), and three on Mar. 9 by L. G. Strolger and R. C. Smith (on behalf of the Supernova Cosmology Project, at the Kitt Peak 0.9-m telescope; cf. IAUC 7125, 7128).

SN 1999 UT α_{2000} δ_{2000} R Offset
1999au Mar. 9 8^{h}58^{m}58.^{s}01 − 7^{o}22′09′′.9 19.2 near center
1999av Mar. 9 10 55 49.70 − 9 20 22.9 18.6 near center
1999aw Mar. 9 11 01 36.37 − 6 06 31.6 17.6  ?
1999ax Mar. 20 14 03 57.92 +15 51 09.2 18.5: 2′′.9 W, 2′′.3 N
1999ay Mar. 21 14 44 43.86 +58 55 42.3 18.0: 2′′.8 W, 2′′.1 S

CCD spectra of SNe 1999au and 1999aw were obtained by A. Goobar, T. Dahlen, and I. Hook on Mar. 16 UT at the Nordic Optical Telescope, from which P. E. Nugent identifies SN 1999au as a type-Ia event (redshift ∼ 0.124) that is likely to be slightly subluminous, resembling the type-Ia SN 1986G (SN 1999au was observed ∼ 1 week before peak brightness, apparently near the center of a faint elliptical galaxy); SN 1999aw is also of type Ia, observed < 1 week after maximum brightness (redshift ∼ 0.04). Based on spectra of SN 1999aw obtained by Strolger and Smith on Mar. 17 at the Cerro Tololo 4-m telescope, M. Phillips suggests that this is likely a luminous, slow-declining type-Ia supernova; no host is apparent in the discovery images. C. Lidman obtained a CCD spectrum of SN 1999av on Mar. 13 at the European Southern Observatory 3.6-m telescope, which shows it to be a type-Ia event at redshift ∼ 0.05, observed 1 week after peak brightness; it is slightly offset from the core of an elliptical galaxy.

SN 1999ax, in the field of Abell 1852 (z = 0.181), was invisible in images obtained on 1998 June 25 (limiting mag R ∼ 22.0). A spectrum (range 330–930 nm) of SN 1999ax obtained by G. Aldering at the Cerro Tololo 4-m reflector on Mar. 22.31 shows a blue continuum with numerous absorption features (widths ∼ 2000 km/s); Nugent and S. Deustua classify the spectrum as that of a very early type-II supernova (yielding z = 0.10, which makes it a foreground object with regard to the galaxy cluster Abell 1852), similar to that of SN 1979C at 7 days past maximum light. Three strong features in the spectrum are present at 380, 475, and 518 nm. SN 1999ay, which is in the field of Abell 1966 (z = 0.151), was confirmed on CCD images taken on Mar. 22 but was invisible in images obtained on 1998 July 27 (limiting mag R ∼ 21.0).

1999 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green